Advertisements
Advertisements
प्रश्न
If `2/11` is the probability of an event, what is the probability of the event ‘not A’.
Advertisements
उत्तर
P(A) = `2/11`
P(not A) = P (A’) = 1 – P(A)
= `1 - 2/11`
= `9/11`
APPEARS IN
संबंधित प्रश्न
A letter is chosen at random from the word ‘ASSASSINATION’. Find the probability that letter is
- a vowel
- an consonant
The probability that a student will pass the final examination in both English and Hindi is 0.5 and the probability of passing neither is 0.1. If the probability of passing the English examination is 0.75, what is the probability of passing the Hindi examination?
In a class of 60 students, 30 opted for NCC, 32 opted for NSS and 24 opted for both NCC and NSS. If one of these students is selected at random, find the probability that
- The student opted for NCC or NSS.
- The student has opted neither NCC nor NSS.
- The student has opted NSS but not NCC.
A die has two faces each with number ‘1’, three faces each with number ‘2’ and one face with number ‘3’. If die is rolled once, determine
- P(2)
- P(1 or 3)
- P(not 3)
Three letters are dictated to three persons and an envelope is addressed to each of them, the letters are inserted into the envelopes at random so that each envelope contains exactly one letter. Find the probability that at least one letter is in its proper envelope.
A dice is thrown. Find the probability of getting a prime number
In a simultaneous throw of a pair of dice, find the probability of getting a doublet of prime numbers
In a simultaneous throw of a pair of dice, find the probability of getting a doublet of odd numbers
In a simultaneous throw of a pair of dice, find the probability of getting neither 9 nor 11 as the sum of the numbers on the faces
In a simultaneous throw of a pair of dice, find the probability of getting a sum more than 6
In a simultaneous throw of a pair of dice, find the probability of getting a sum less than 7
In a simultaneous throw of a pair of dice, find the probability of getting a sum more than 7
In a simultaneous throw of a pair of dice, find the probability of getting odd number on the first and 6 on the second
In a simultaneous throw of a pair of dice, find the probability of getting a total of 9 or 11
In a simultaneous throw of a pair of dice, find the probability of getting a total greater than 8.
Three coins are tossed together. Find the probability of getting exactly two heads
Three coins are tossed together. Find the probability of getting at least two heads
Three coins are tossed together. Find the probability of getting at least one head and one tail.
A box contains 10 red marbles, 20 blue marbles and 30 green marbles. 5 marbles are drawn at random. From the box, what is the probability that all are blue?
A box contains 6 red marbles numbered 1 through 6 and 4 white marbles numbered from 12 through 15. Find the probability that a marble drawn is white .
A box contains 6 red marbles numbered 1 through 6 and 4 white marbles numbered from 12 through 15. Find the probability that a marble drawn is white and odd numbered .
A box contains 6 red marbles numbered 1 through 6 and 4 white marbles numbered from 12 through 15. Find the probability that a marble drawn is red or even numbered.
Fill in the blank in the table:
| P (A) | P (B) | P (A ∩ B) | P(A∪ B) |
| 0.35 | .... | 0.25 | 0.6 |
If A and B are two events associated with a random experiment such that P(A) = 0.3, P (B) = 0.4 and P (A ∪ B) = 0.5, find P (A ∩ B).
If A and B are two events associated with a random experiment such that
P(A) = 0.5, P(B) = 0.3 and P (A ∩ B) = 0.2, find P (A ∪ B).
If A and B are two events associated with a random experiment such that
P (A ∪ B) = 0.8, P (A ∩ B) = 0.3 and P \[(\bar{A} )\]= 0.5, find P(B).
There are three events A, B, C one of which must and only one can happen, the odds are 8 to 3 against A, 5 to 2 against B, find the odds against C
A card is drawn from a deck of 52 cards. Find the probability of getting an ace or a spade card.
An urn contains 9 balls two of which are red, three blue and four black. Three balls are drawn at random. The probability that they are of the same colour is
A box contains 6 nails and 10 nuts. Half of the nails and half of the nuts are rusted. If one item is chosen at random, the probability that it is rusted or is a nail is
Three numbers are chosen from 1 to 20. The probability that they are not consecutive is
In a certain lottery 10,000 tickets are sold and ten equal prizes are awarded. What is the probability of not getting a prize if you buy two tickets.
